# Superluminal Wave Activation at Relativistic Magnetized Shocks

### v1

@@ IV.3 Astrophysical Implications: Fast Radio Bursts
Plotnikov2019 and Sironi2021 analyzed kinetic plasma simulations of shock maser emission. This coherent process generates radiation, possibly at FRB frequencies. However, its efficiency decreases for increasing magnetization and depends on upstream conditions like temperature (Babul2020; Vanthieghem2025a). Future parameter scans of the wave activation mechanism discussed in this work will evaluate the simultaneous occurrence of synchrotron maser emission and mode activation. In the transition from mild ($\sigma\approx 1$) to strong magnetizations, wave activation at the shock may alleviate the efficiency constraints of precursor emission (Thompson2022).
@@
This letter is a first-principles proof of concept for a mechanism that activates non-propagating Alfvénic perturbations at relativistic magnetized shocks (Thompson2022). Small-scale fluctuations at scales above a critical wavenumber pass through the shock, generating propagating superluminal modes downstream. Depending on the flow and plasma properties, the activated waves could reach GHz frequencies as a possible source of FRBs.
@@
We thank Ludwig Böss, Hayk Hakobyan, Amir Levinson, Alexander Philippov, Siddhant Solanki, Anatoly Spitkovsky, and Chris Thompson for valuable discussions. J.F.M. acknowledges support from NSF grant AST-2508744. L.S. acknowledges support from the DOE Early Career Award DE-SC0023015, NASA ATP 80NSSC24K1826, NSF AST-2307202, and the Simons Foundation (MP-SCMPS-00001470). This work was facilitated by the Multimessenger Plasma Physics Center (NSF PHY-2206609) and supported in part by NSF PHY-2309135 to the Kavli Institute for Theoretical Physics. We acknowledge GitHub Copilot for coding assistance, and generative AI tools for figure design. Simulations were performed on the Ginsburg (Columbia University), Rusty (Flatiron Institute), and Discovery (Dartmouth College) clusters.
@@ A.1 Relevant frame transformations and identities
Equation (3) provides the dispersion relation for superluminal O-modes in the plasma rest frame. We derive the corresponding dispersion relation in the shock frame (s) by substituting primed quantities into Equation (3):
@@
Applying the identity $1-\beta^{2}=1/\gamma^{2}$, we obtain the O-mode dispersion in the shock-front frame:
@@
Quantity Downstream (d) Shock-Front (s) Upstream (u) Notes
@@ A.2 Shock structure from homogeneous and strongly magnetized shell collisions
The astrophysical interpretation of the mode conversion mechanism evaluated in this letter relies on the shock structure induced during collisions of uniform and highly magnetized shells discussed by Thompson2022 and, more generally, by Beloborodov2020. We consider a structure of reverse shock (rs), contact discontinuity (c), and forward shock (fs), propagating through a relativistic wind with an upstream (u) velocity $\gamma_{u|\ast}$ and a downstream (d) velocity $\gamma_{d|\ast}$. The specific solution discussed here has $\gamma_{{\rm rs}|\ast}\ll\gamma_{{\rm c}|\ast}\ll\gamma_{{\rm fs}|\ast}$. The MHD jump conditions imply continuous tangential electric fields in the shock-front-frame:
@@ A.3 Jump Conditions
Continuity, momentum, and energy conservation in ideal relativistic MHD provide shock jump conditions. In the shock-front frame, we find (e.g., Lemoine2016):
@@
Here, double brackets $\dots$]] denote the difference of a quantity between immediate downstream and upstream. The normal vector in the shock-front frame is $n_{\alpha}=[0,1,0,0]$. $J^{\alpha}=\rho u^{\alpha}$ is the four-current in terms of the flow density $\rho$ and shock four-velocity $u^{\alpha}=[\gamma_{\rm s},\gamma_{\rm s}\beta_{\rm s},0,0]$, $T^{\alpha\beta}$ refers to the energy-momentum tensor. $^{*}F^{\alpha\beta}$ is the (dual) electromagnetic strength tensor, commonly expressed in terms of the shock velocity and magnetic field four-vector $^{*}F^{\alpha\beta}=u^{\alpha}b^{\beta}-u^{\beta}b^{\alpha}$. For perpendicular shocks (Section III.1) we find the following four conserved quantities across the shock:
@@ Appendix B Mode activation for spectra of seed waves
Figure 6: Same as Figure 4, but for a spectrum of seed perturbations with $s=1$ sampled with $N=100$ modes. We show the field components (panel a), a zoom of the wave EM fields (panel b), and a Fourier spectrum of the wave magnetic field in different regions (panel c). The Fourier spectrum distinguishes the immediate and far downstream (c, left panel), and the seed perturbations upstream of the shock (c, right panel). The red dashed line estimates the cutoff (Equation 8) for downstream scales, as outlined in Appendix B. The plasma frequency cutoff acts like a high-pass filter for waves in the far downstream.
We extend the analysis of Section III.3 to a spectrum of seed waves sampled from an energy distribution with decay index $s=1$. Figure 6 shows the corresponding wave activation dynamics. As outlined in Figure 4, seed waves at scales above the cutoff (Equation 8) are activated and propagate far downstream of the shock. From Equation (Equation 8) and a suitable length contraction of the wave scale, we estimate the downstream cutoff as

### v2

@@ IV.3 Astrophysical Implications: Fast Radio Bursts
Plotnikov and Sironi (2019) and Sironi et al. (2021) analyzed kinetic plasma simulations of shock maser emission. This coherent process generates radiation, possibly at FRB frequencies. However, its efficiency decreases for increasing magnetization and depends on upstream conditions like temperature (Babul and Sironi, 2020; Vanthieghem and Levinson, 2025b). Future parameter scans of the wave activation mechanism discussed in this work will evaluate the simultaneous occurrence of synchrotron maser emission and mode activation. In the transition from mild ($\sigma\approx 1$) to strong magnetizations, wave activation at the shock may alleviate the efficiency constraints of precursor emission (Thompson, 2022).
@@
This letter is a first-principles proof of concept for a mechanism that activates nonpropagating Alfvénic perturbations at relativistic magnetized shocks (Thompson, 2022). Small-scale fluctuations at scales above a critical wavenumber pass through the shock, generating propagating superluminal modes downstream. Depending on the flow and plasma properties, the activated waves could reach GHz frequencies as a possible source of FRBs.
@@
We are grateful for the anonymous referee’s helpful comments. We thank Ludwig Böss, Hayk Hakobyan, Amir Levinson, Alexander Philippov, Siddhant Solanki, Anatoly Spitkovsky, and Christopher Thompson for valuable discussions. We thank Nicole Wotring for supporting the science communication efforts related to this work by producing animations based on the simulation data. L.E. and N.W. are grateful for the support by Dartmouth College’s SURFD undergraduate research programs. J.F.M. acknowledges support from NSF grant AST-2508744. L.S. acknowledges support from the DOE Early Career Award DE-SC0023015, NASA ATP 80NSSC24K1826, NSF AST-2307202, and the Simons Foundation (MP-SCMPS-00001470). This work was facilitated by the Multimessenger Plasma Physics Center (NSF PHY-2206609) and supported in part by NSF PHY-2309135 to the Kavli Institute for Theoretical Physics. We acknowledge GitHub Copilot for coding assistance, and generative AI tools for figure design. Simulations were performed on the Ginsburg (Columbia University), Rusty (Flatiron Institute), and Discovery (Dartmouth College) clusters. This work was granted access to the HPC resources of TGCC/CCRT under the allocation A0180415130 made by GENCI.
@@ A.1 Relevant Frame Transformations and Identities
Equation (3) provides the dispersion relation for superluminal O-modes in the plasma rest-frame. We derive the corresponding dispersion relation in the shock frame (s) by substituting primed quantities into Equation (3):
@@
Applying the identity $1-\beta^{2}=1/\gamma^{2}$, we obtain the O-mode dispersion in the shock front frame:
@@
Quantity Downstream (d) Shock front (s) Upstream (u) Notes
@@ A.2 Shock structure from homogeneous and strongly magnetized shell collisions
The astrophysical interpretation of the mode conversion mechanism evaluated in this letter relies on the shock structure induced during collisions of uniform and highly magnetized shells discussed by Thompson (2022, Section 6) and, more generally, by Beloborodov (2020). We consider a structure of reverse shock (rs), contact discontinuity (c), and forward shock (fs), propagating through a relativistic wind with an upstream (u) velocity $\gamma_{u|\ast}$ and a downstream (d) velocity $\gamma_{d|\ast}$. The specific solution discussed here has $\gamma_{{\rm rs}|\ast}\ll\gamma_{{\rm c}|\ast}\ll\gamma_{{\rm fs}|\ast}$. The MHD jump conditions imply continuous tangential electric fields in the shock-front-frame:
@@ A.3 Jump Conditions
Continuity, momentum, and energy conservation in ideal relativistic MHD provide shock jump conditions. In the shock front frame, we find (e.g., Lemoine et al., 2016):
@@
Here, double brackets $\dots$]] denote the difference of a quantity between immediate downstream and upstream. The normal vector in the shock front frame is $n_{\alpha}=[0,1,0,0]$. $J^{\alpha}=\rho u^{\alpha}$ is the four-current in terms of the flow density $\rho$ and shock four-velocity $u^{\alpha}=[\gamma_{\rm s},\gamma_{\rm s}\beta_{\rm s},0,0]$, $T^{\alpha\beta}$ refers to the energy-momentum tensor. $^{*}F^{\alpha\beta}$ is the (dual) electromagnetic strength tensor, commonly expressed in terms of the shock velocity and magnetic field four-vector $^{*}F^{\alpha\beta}=u^{\alpha}b^{\beta}-u^{\beta}b^{\alpha}$. For perpendicular shocks (Section III.1) we find the following four conserved quantities across the shock:
@@ Appendix B Mode activation for spectra of seed waves
Figure 6: Same as Figure 4, but for a spectrum of seed perturbations with $s=1$ sampled with $N=100$ modes. We show the field components (panel a), a zoom of the wave EM fields (panel b), and a Fourier spectrum of the wave magnetic field in different regions (panel c). The Fourier spectrum distinguishes the immediate and far downstream (c, left panel), and the seed perturbations upstream of the shock (c, right panel). The red dashed line estimates the cutoff (cold plasma estimate, Equation 8) for downstream scales, as outlined in Appendix B. The plasma frequency cutoff acts like a high-pass filter for waves in the far downstream.
We extend the analysis of Section III.3 to a spectrum of seed waves sampled from an energy distribution with decay index $s=1$. Figure 6 shows the corresponding wave activation dynamics. As outlined in Figure 4, seed waves at scales above the cutoff (Equation 8) are activated and propagate far downstream of the shock. From Equation (Equation 8) and a suitable length contraction of the wave scale, we estimate the downstream cutoff in the cold plasma limit as